OpenStudy (anonymous):

Find the solution of 3 times the square root of the quantity of x plus 6 equals negative 12, and determine if it is an extraneous solution

OpenStudy (anonymous):

\[3\sqrt{x+6}=-12\]

TheSmartOne (thesmartone):

Divide 3 on both sides first.

OpenStudy (anonymous):

\[\sqrt{x+6}=-4\] Then I multiply by -1 on each side?

TheSmartOne (thesmartone):

No, we square both sides.

OpenStudy (anonymous):

oh that's right this isn't an I equation so itll be x=10 after squaring and subtracting 6

TheSmartOne (thesmartone):

and now, plug back in x = 10 when you plug back in x = 10, does it work?

OpenStudy (anonymous):

no it doesn't

TheSmartOne (thesmartone):

so that means it is an extraneous solution

OpenStudy (anonymous):

oh that was pretty easy thought I got the wrong x value thanks again :)